HW 17, Problem 3.

Compute the characteristic values and corresponding characteristic vectors of the given matrix. Write the vectors in their general form and give a specific numerical example. Also prove that if A is the given original Matrix and D, is its diagonization matrix then A and D are similar.

4 1 6 6 3 2 -5 -2
Answer #1

Given matrix is 4 6 6 A = 1 3 -5 2. - 2 The characteristic equation of A is IA-d Il 0 lie. lu-d 6 6. 3-d 2 -) - 5 -2-d TT 2-5R2 R2 - Ri 2 2 TE Rs - R3+R, -39-2, = 0, 2, + 24, + 22, = 0 2 =-34, and x + 2y + 2 (-34, ) = 0 2,=44, Hence, 4y, g, Z y (-34,1 3 3 R3 R3-R2 -2 -1 o 22 0 -242-22 and Ka+ 3y2 + 3 22=0 x2 + 3y2 + 3(-242) = 0 1 22=-242 =) X 2 = 342 Hence, X2 Yz 22 342 Y2